# Enhancing reliability in prediction intervals using point forecasters: Heteroscedastic Quantile Regression and Width-Adaptive Conformal Inference

### v1

@@ 5.1 Electricity Price Forecasting (EPF)
The results for the EPF data are shown in Table 1. As the hourly quantile regression models performed considerably worse compared to the consideration of a single time series, only the results for the latter situation are shown. The results of the QRA and HQR models by hours are shown in B.
For every $\alpha$, it is seen that the HQR model performs better than the QRA model. In particular, higher coverage values are obtained with smaller intervals, as shown by the mean or median interval length. This results in noticeable differences in the Winkler score. The differences between the models are not reduced by applying conformal procedures, they remain at the same level. Thus, the HQR model is a superior option to the QRA. The standard CQR procedure contributes nothing in particular, it does not improve in any case the results of quantile regression alone. For $\alpha=0.01$ or $\alpha=0.05$ the intervals are worse. For $\alpha=0.10$ or $\alpha=0.20$ the intervals are similar. This could be due to the absence of the exchangeability condition in the series. With respect to the ACI method, there is a tendency to undercover compared to the rest of the methods. In particular, in the extreme case of $\alpha=0.01$, the coverage is well below the 99% desired and this undercoverage is not rewarded by making very efficient intervals, as the Winkler score is very large. Even so, for the rest of the cases, the Winkler score is the best or very similar to the best. The WACI method always obtains valid intervals and the Winkler score levels are always comparable to the best. The performance on the extreme case at 99% is very remarkable, where it is the only conformal method that keeps up with the HQR with respect to the Winkler score and with almost exact coverage. From now on, only the HQR and QRA models are considered, as well as the different conformalizations of the former due to its superior performance.
With regard to the distribution of the length of the intervals, Figures 9, 10, 11 and 12 summarise this aspect.
Method Obj. Coverage = $99\%\,(\alpha=0.01)$ Obj. Coverage = $95\%\,(\alpha=0.05)$
Emp. Cov. Mean I.L. Median I.L. WS Emp. Cov. Mean I.L. Median I.L. WS
HQR 99,11 109,32 110,04 126,32 96,12 74,59 75,90 91,00
QRA 98,88 114,80 116,35 139,39 95,98 78,39 81,23 96,34
CQR-HQR 99,26 118,02 115,45 130,38 96,90 79,67 81,81 92,45
CQR-QRA 99,48 133,39 134,55 143,32 96,62 84,14 86,81 97,56
ACI-HQR 97,72 89,63 88,11 133,02 94,91 72,92 72,56 91,33
ACI-QRA 97,89 100,00 99,28 144,94 94,79 77,16 78,44 96,42
WACI-HQR 99,04 111,20 107,29 127,57 96,58 78,38 80,59 92,43
WACI-QRA 99,27 124,38 124,57 137,41 96,63 82,74 86,88 97,07
Method Obj. Coverage = $90\%\,(\alpha=0.10)$ Obj. Coverage = $80\%\,(\alpha=0.20)$
Emp. Cov. Mean I.L. Median I.L. WS Emp. Cov. Mean I.L. Median I.L. WS
HQR 92,52 58,12 59,15 74,87 84,15 42,11 43,34 59,36
QRA 92,39 59,52 61,67 77,94 84,27 42,91 44,66 60,88
CQR-HQR 92,39 59,05 61,02 74,92 82,92 42,03 43,80 59,32
CQR-QRA 92,45 61,42 64,01 78,21 83,27 43,02 45,01 60,73
ACI-HQR 89,75 54,95 54,22 74,23 79,61 39,15 38,20 58,90
ACI-QRA 89,82 56,80 56,22 77,37 79,59 39,69 39,39 60,30
WACI-HQR 92,00 58,06 60,15 74,64 82,16 41,01 42,60 59,05
WACI-QRA 92,25 61,21 64,90 77,64 82,92 42,23 44,55 60,03
Table 1: Results in the EPF example.
Figure 9: Distribution of the interval length for $\alpha=0.01$ in the EPF dataset.
Figure 10: Distribution of the interval length for $\alpha=0.05$ in the EPF dataset.
Figure 11: Distribution of the interval length for $\alpha=0.10$ in the EPF dataset.
Figure 12: Distribution of the interval length for $\alpha=0.20$ in the EPF dataset.
The first point to note is that the HQR method and QRA show relatively similar behaviour for every $\alpha$, but with considerably smaller intervals across the whole distribution for the former. This explains why the observed Winkler score was so much better (Table 1). A clear difference can also be observed in the upper tail of the distribution, where situations of higher uncertainty are clearly differentiated from the rest in the case of the HQR. Interestingly, the ACI method shows the smallest intervals throughout the distribution except for the upper quantiles, where the jump is noticeable and it becomes the method with the largest intervals. This happens for all $\alpha$, although for $\alpha=0.01$ the ACI method does not obtain valid intervals. Take into account that this property could not be observed with just the mean or the median interval length. The comparison between the HQR method and its WACI variation is of interest: WACI prefers to use smaller intervals in the lower tail, larger intervals in the upper tail and shows similar behaviour in situations close to the median. This is one of the key methodological differences with ACI, as a different adaptation can be seen depending on the interval length. An important aspect to consider is that the intervals produced through conformal procedures are more diverse. Although the validity and efficiency metrics are similar to not applying them, we can differentiate situations more easily when conformal methods are applied.
As has been explained several times in the paper, although the ranges are more varied for HQR and its conformal modifications than for QRA, this variation must be linked to the difficulty of the prediction. Figures 13, 14, 15 and 16 discuss this quality.
The construction of the HQR model has focused on this characteristic. Every $\alpha$ show the same properties: the HQR model captures this behaviour very well and the length of the interval is correctly related to the error being made. However, the QRA model does not capture the uncertainty well, except perhaps for the case $\alpha=0.05$. Thus, it can be said that our intervals are informative, whereas the intervals produced by QRA can lead to errors in decision making by presenting inaccuracies (like saying that there is little uncertainty in situations of large error) in this regard. Furthermore, in this case also the conformal methods are slightly superior to HQR. Although all of them, except QRA, produce the desired increasing relationship, the conformal procedures manage to avoid even slight deviations. For example, look at the difference of groups $(0.4,0.5)$ and $(0.5,0.6)$ with groups $(0.6,0.7)$ and $(0.7,0.8)$ in Figures 13, 15 and 16.
Figure 13: MAE per group of observations for the EPF dataset, where the groups are formed according to the length of the intervals for $\alpha=0.01$
Figure 14: MAE per group of observations for the EPF dataset, where the groups are formed according to the length of the intervals for $\alpha=0.05$
Figure 15: MAE per group of observations for the EPF dataset, where the groups are formed according to the length of the intervals for $\alpha=0.10$
Figure 16: MAE per group of observations for the EPF dataset, where the groups are formed according to the length of the intervals for $\alpha=0.20$
Finally, it will be checked that, when considering a set of intervals, the coverage does not depend on the length of the interval. That is, on the degree of difficulty of forecasting. For this purpose, the charts and the measures described in Section 4.2 are plotted/computed. For ease of comprehension, only the graphs considering 5% observations per group are shown. These are given in Figures 17, 18, 19 and 20. Table 2 shows the results of the coverage deviations for all cases.
For $\alpha=0.01$ there is by itself not much scope for overcoverage and nothing particular is observed for this level in general, where all methods behave at acceptable levels even at their maximum deviation. However, for the rest of the $\alpha$ values the performance of the ACI method is of concern. The tendency of the method to undercoverage has been observed. However, it can be found here that this undercoverage is also related to the smaller intervals for the most part. If we look at the 5% groups of observations, the deviation from the target coverage is as high as 25%, which is unacceptable. In fact, an overconfidence bias is observed in the 10% of observations with smaller intervals. The WACI method corrects ACI by mitigating this bias. In fact, overall, it is either the best, or at similar levels compared to the best in terms of deviations by having the good properties of ACI, without the problems it presents in this regard. The quantile regression models themselves in fact appear to be well calibrated on their own in this respect, with no noticeable biases observed. One problem that all methods seem to present is that for observations with larger intervals there is underconfidence. In that case the smaller intervals would be preferred, which are associated with the QRA method (Figures 9 to 12), but this method does not provide fully informative intervals (Figures 13 to 16), so the HQR model would be the best choice for such cases. Moreover, if in Table 2 one looks at the deviations of each method for the 5% groups, the WACI method is not the best by a noticeable margin, however the few observations per group may distort the results somewhat, as only one year (365 days per hour) is being analysed and the fact that the conformalization is done by hour and by interval length makes the number of samples for each interval very small. Analizing the 10% groups where the number of samples per interval is big enough, the WACI method is clearly the one with smallest deviations except in the case $\alpha=0.05$, where it is in line with the best. Notably, improvement of maximum deviations with respect to HQR are around 20% for $\alpha=0.10$ and $\alpha=0.20$.
Figure 17: Empirical coverage per group of observations for the EPF dataset, where the groups are formed according to the length of the intervals for $\alpha=0.01$
Figure 18: Empirical coverage per group of observations for the EPF dataset, where the groups are formed according to the length of the intervals for $\alpha=0.05$
Figure 19: Empirical coverage per group of observations for the EPF dataset, where the groups are formed according to the length of the intervals for $\alpha=0.10$
Figure 20: Empirical coverage per group of observations for the EPF dataset, where the groups are formed according to the length of the intervals for $\alpha=0.20$
$\alpha$ 0.01
% obs. per group 5% 10%
Method Mean Deviation Max Deviation Mean Deviation Max Deviation
HQR 0.61 2.64 0.56 1.16
QRA 0.76 2.64 0.74 1.62
CQR-HQR 0.60 1.05 0.50 0.93
ACI-HQR 1.36 3.77 1.29 2.87
WACI-HQR 0.56 1.28 0.46 0.93
$\alpha$ 0.05
% obs. per group 5% 10%
Method Mean Deviation Max Deviation Mean Deviation Max Deviation
HQR 1.83 4.09 1.60 3.18
QRA 1.69 3.41 1.58 3.41
CQR-HQR 2.20 4.32 2.00 3.18
ACI-HQR 2.17 13.18 2.03 7.86
WACI-HQR 1.93 4.09 1.63 3.52
$\alpha$ 0.10
% obs. per group 5% 10%
Method Mean Deviation Max Deviation Mean Deviation Max Deviation
HQR 3.46 7.49 2.98 5.79
QRA 3.27 7.73 2.77 7.16
CQR-HQR 3.42 9.32 3.11 4.99
ACI-HQR 3.07 20.91 3.07 13.21
WACI-HQR 2.76 7.50 2.69 4.88
$\alpha$ 0.20
% obs. per group 5% 10%
Method Mean Deviation Max Deviation Mean Deviation Max Deviation
HQR 5.06 12.48 4.50 10.32
QRA 4.79 13.85 4.26 11.69
CQR-HQR 4.96 18.18 4.91 9.92
ACI-HQR 4.52 25.00 3.95 17.20
WACI-HQR 4.04 13.41 3.97 7.99
Table 2: Mean and maximum deviation of each model from the objective coverage by interval length for the EPF dataset.
As a final analysis, the effect of hourly conformalization has been studied. The empirical hourly coverage is shown in Figures 21, 22, 23 and 24.
Figure 21: Coverage by hour for $\alpha=0.01$ in the EPF dataset.
Figure 22: Coverage by hour for $\alpha=0.05$ in the EPF dataset.
Figure 23: Coverage by hour for $\alpha=0.10$ in the EPF dataset.
Figure 24: Coverage by hour for $\alpha=0.20$ in the EPF dataset.
In these figures, the tendency of the ACI method to undercoverage can be better appreciated, especially in the case of target coverage equal to 99%. Even so, this method presents a very good characteristic in the uniformity of the distribution of coverage between hours. No other method achieves such good results in this sense. For further understading, the average deviation of each hour from the target coverage, the maximum deviation and the standard deviation of the deviations (as a measure of uniformity) have been computed. The results are shown in Tables 3 and 4.
$\alpha$ 0.01 0.05
Method
Mean
deviation
Max
deviation
Std.
Mean
deviation
Max
deviation
Std.
HQR 0.42 1 0.48 1.23 2.54 0.91
QRA 0.33 1.19 0.42 1.07 2.81 0.86
CQR-HQR 0.37 0.73 0.34 1.90 3.63 0.70
ACI-HQR 1.28 2 0.32 0.19 0.74 0.25
WACI-HQR 0.30 0.91 0.35 1.58 2.27 0.55
Table 3: Mean deviation, maximum deviation and standard deviation of the deviations of the coverage by hours in the EPF dataset. $\alpha=0.01$ and $\alpha=0.05$
$\alpha$ 0.10 0.20
Method
Mean
deviation
Max
deviation
Std.
Mean
deviation
Max
deviation
Std.
HQR 2.61 5.08 1.65 4.15 10.16 2.45
QRA 2.39 4.26 1.36 4.31 9.34 2.73
CQR-HQR 2.39 4.26 1.05 2.94 7.70 1.95
ACI-HQR 0.29 0.66 0.24 0.48 1.04 0.39
WACI-HQR 2 3.72 0.90 2.16 4.15 1.04
Table 4: Mean deviation, maximum deviation and standard deviation of the deviations of the coverage by hours in the EPF dataset. $\alpha=0.10$ and $\alpha=0.20$
The positive comments on the ACI method are confirmed by the results. However, the WACI method shows exceptional performance in the case of the most extreme quantiles ($\alpha=0.01$) and shows more than competent results at all other levels. In fact, the conformalization step adds value in this respect, as even the CQR method improves results over the quantile regression alone.
As the conformalization is being done on an hourly basis, what we are really looking at here is the marginal coverage of conformalization for each hour. Both ACI and WACI present theoretical guarantees that with infinite timesteps, this coverage has the desired level. However, it can be seen from these graphs that ACI achieves these results faster than WACI. This is because ACI only makes modifications based on the last observation without discerning by interval length. In the case of WACI, you need many observations of all interval sizes considered, which is unlikely in the studied period (365 days). Thus WACI, despite showing good results, does not achieve the same level of performance as ACI in this respect. The performance should improve and be in line with the ACI method with enough observations.
In conclusion, HQR has shown clear superiority to QRA in all aspects. Particularly, a big difference is observed in a correct quantification of the uncertainty by showing a relationship between the difficulty of the prediction and the length of the interval. The different conformalizations do not affect most of the characteristics and comparing the classical metrics, both ACI and WACI could be considered as the best models indistinctly. However, a clear bias can be seen in ACI by showing situations of overconfidence for the simplest predictions. WACI corrects this bias without considerably worsening any other characteristics. Everything considered, the WACI-HQR model would be the best in this example.
5.2 Wind Power Forecasting (WPF)
The results associated with this example are given in detail in Table 5. This example shows a completely different behaviour from the previous one. In the case of EPF, there was a very noticeable gap between the QRA and HQR models. In this case, the two models show practically the same results for all levels of miscoverage targets. The same effect can be seen with the conformalizations. This is because for a large part of the time, the standard deviation feature is probably not significant in HQR. If that is the case, you really have two models that use only one estimator of the mean of the distribution of interest. In the HQR model all predictions are given equal weight and thus a single coefficient is used in the model. In the QRA model it is estimated as a weighted mean, resulting in $M$ explanatory variables. This is discussed in A. However, there is a noticeable improvement in the adaptive conformal procedures, which in the previous case was not directly evident, but rather through the other characteristics studied. Between the ACI and WACI conformalizations, no major differences can be observed in this case. In the previous example ACI gave considerably smaller intervals, but penalised coverage significantly, whereas WACI aimed for efficiency within validity. In any case, the Winkler score values were similar. In this example, the Winkler score values are still comparable, but with both methods showing the same type of behaviour. Only slight differences are observed for $\alpha=0.01$, where the WACI methodology shows the better results. For this same value of objective miscoverage, the HQR model shows clear superiority with respect to QRA, as was also the case in the previous example. It can also be seen again that the CQR conformalization does not bring any improvement, as expected. In view of these results, only the remaining characteristics for the ACI and WACI models are analysed.
With regard to the distribution of the length of the intervals, Figures 25, 26, 27 and 28 summarise this aspect. The differences between ACI and WACI are much smaller in this case, regardless of the level of $\alpha$. This was expected as it was already noted when discussing Table 5. The same behaviour is observed for every $\alpha$, except for $\alpha=0.01$, where the length of the intervals is smaller for the hole distribution for the ACI methodology, but penalising the coverage. For the rest of $\alpha$, WACI produces slightly larger intervals than ACI for about 40% of the smallest intervals and smaller for the top 40%, except for just the most extreme intervals, while the behaviour in the central area around the median is similar. There is no difference between QRA and HQR in this matter.
Method Obj. Coverage = $99\%\,(\alpha=0.01)$ Obj. Coverage = $95\%\,(\alpha=0.05)$
Emp. Cov. Mean I.L. Median I.L. WS Emp. Cov. Mean I.L. Median I.L. WS
HQR 98,40 4,09 3,88 5,31 94,15 2,97 2,88 3,95
QRA 98,12 4,03 3,86 5,45 93,70 2,94 2,85 3,95
CQR-HQR 98,53 4,34 4,07 5,35 94,73 3,10 2,92 3,90
CQR-QRA 98,42 4,35 4,06 5,32 94,85 3,08 2,94 3,92
ACI-HQR 96,61 3,51 3,33 5,46 93,79 2,80 2,67 3,54
ACI-QRA 96,68 3,50 3,29 5,50 93,63 2,79 2,68 3,53
WACI-HQR 97,11 3,63 3,40 5,33 94,09 2,85 2,71 3,56
WACI-QRA 97,07 3,66 3,40 5,43 94,02 2,84 2,74 3,59
Method Obj. Coverage = $90\%\,(\alpha=0.10)$ Obj. Coverage = $80\%\,(\alpha=0.20)$
Emp. Cov. Mean I.L. Median I.L. WS Emp. Cov. Mean IL Median IL WS
HQR 88,54 2,42 2,35 3,37 78,45 1,81 1,77 2,75
QRA 88,70 2,39 2,33 3,36 78,68 1,81 1,77 2,75
CQR-HQR 90,02 2,54 2,45 3,34 79,93 1,89 1,84 2,73
CQR-QRA 89,83 2,50 2,43 3,33 80,10 1,87 1,84 2,73
ACI-HQR 89,35 2,41 2,31 3,07 79,70 1,88 1,82 2,58
ACI-QRA 89,28 2,38 2,32 3,06 79,75 1,86 1,81 2,57
WACI-HQR 89,51 2,39 2,31 3,08 79,92 1,87 1,79 2,58
WACI-QRA 89,51 2,37 2,30 3,07 79,95 1,86 1,79 2,57
Table 5: Results in the WPF example.
Figure 25: Distribution of the interval length for $\alpha=0.01$ in the WPF dataset.
Figure 26: Distribution of the interval length for $\alpha=0.05$ in the WPF dataset.
Figure 27: Distribution of the interval length for $\alpha=0.10$ in the WPF dataset.
Figure 28: Distribution of the interval length for $\alpha=0.20$ in the WPF dataset.
Figures 29, 30, 31 and 32 discuss the relationship between the interval length and the difficulty of the prediction.
Figure 29: MAE per group of observations for the WPF dataset, where the groups are formed according to the length of the intervals for $\alpha=0.01$
Figure 30: MAE per group of observations for the WPF dataset, where the groups are formed according to the length of the intervals for $\alpha=0.05$
Figure 31: MAE per group of observations for the WPF dataset, where the groups are formed according to the length of the intervals for $\alpha=0.10$
Figure 32: MAE per group of observations for the WPF dataset, where the groups are formed according to the length of the intervals for $\alpha=0.20$
Both QRA and HQR are able to capture well the difficulty of the prediction with the length of the interval. This is consistent with the fact that the standard deviation of the models is not a significant variable. The models do not need such a variable to correctly quantify uncertainty for this example, hence QRA is able to perform this function well. We have studied this property when the coefficients associated to the estimation of the standard deviation between the predictors, $\lambda_{2}(\frac{\alpha}{2})$ and $\lambda_{2}(1-\frac{\alpha}{2})$, are significant or not. The results are described in A.
Regarding the coverage by interval length, Figures 33, 34, 35 and 36 are analysed. As in the case of the EPF, only the plots associated with the groups with 5% observations are presented, although Table 6 shows the results for the different configurations. Table 6 shows how the effect of WACI is positive. Although overall, the maximum deviation from the target coverage is not as large, WACI achieves better results for the mean and maximum deviation for every $\alpha$ condsidered. In particular, it should be noted that the maximum deviation of WACI when considering groups of 10% of observations is very small. Moreover, from the Figures 33 to 36, it can be seen that these improvements come mainly from correcting the overconfidence of ACI in the smallest and largest intervals, which corresponds to the effect observed in Figures 29 to 32. Furthermore, ignoring the particular behavior of the extreme case, WACI seems to achieve similar results for each group. In other words, the distribution of coverage according to the groups considered is more uniform for WACI than for ACI, as shown in Figures 34, 35 and 36.
To conclude the discussion on this example, it has been observed that, overall, there are no notable differences between QRA and HQR, although they may exist in specific time periods in a favorable way for HQR (A). Standard CQR conformalization still does not provide great results on this problem. Although on classical measures, ACI and WACI obtain essentially the same results, WACI again corrects for the overconfidence biases exhibited by ACI, again making the WACI procedure more appropriate.
Figure 33: Empirical coverage per group of observations for the WPF dataset, where the groups are formed according to the length of the intervals for $\alpha=0.01$
Figure 34: Empirical coverage per group of observations for the WPF dataset, where the groups are formed according to the length of the intervals for $\alpha=0.05$
Figure 35: Empirical coverage per group of observations for the WPF dataset, where the groups are formed according to the length of the intervals for $\alpha=0.10$
Figure 36: Empirical coverage per group of observations for the WPF dataset, where the groups are formed according to the length of the intervals for $\alpha=0.20$
$\alpha$ 0.01
% obs. per group 5% 10%
Method Mean Deviation Max Deviation Mean Deviation Max Deviation
ACI-QRA 2.34 4,71 2.32 3.91
WACI-QRA 1.99 5.85 1.93 3.45
ACI-HQR 2.42 5.39 2.39 4.37
WACI-HQR 1.91 4.94 1.89 4.71
$\alpha$ 0.05
% obs. per group 5% 10%
Method Mean Deviation Max Deviation Mean Deviation Max Deviation
ACI-QRA 1.55 4.13 1.48 3.56
WACI-QRA 1.11 3.45 0.98 1.74
ACI-HQR 1.90 5.27 1.77 3.56
WACI-HQR 1.13 2.99 0.91 1.96
$\alpha$ 0.10
% obs. per group 5% 10%
Method Mean Deviation Max Deviation Mean Deviation Max Deviation
ACI-QRA 1,55 4,13 1,48 3,56
WACI-QRA 1,11 3,45 0,98 1,74
ACI-HQR 1,90 5,27 1,77 3,56
WACI-HQR 1,13 2,99 0,91 1,96
$\alpha$ 0.20
% obs. per group 5% 10%
Method Mean Deviation Max Deviation Mean Deviation Max Deviation
ACI-QRA 2.63 6.76 2.10 5.27
WACI-QRA 2.10 5.39 1.83 5.05
ACI-HQR 2.89 7.40 2.45 6.37
WACI-HQR 2.16 4.66 1.93 3.22
Table 6: Mean and maximum deviation of each model from the objective coverage by interval length for the WPF dataset.

### v2

@@ 4.3 Electricity Price Forecasting (EPF)
Electricity Price Forecasting serves as an ideal example for evaluating different techniques in the area of probabilistic forecasting. Producing prediction intervals in this context is of significant interest due to the intricate dynamics of electricity markets, which are increasingly influenced by stochastic factors, such as renewable energy integration. These factors directly impact the strategies of market participants and the formulation of their bids. As a result, the area of probabilistic forecasting has gained traction, with numerous studies using the electricity market as a benchmark for evaluating their methodologies. For instance, Wisniewski et al. (2020); Kath and Ziel (2021); Zaffran et al. (2022).
Figure 8: (a) One week example of the point forecasters for the EPF example and (b) time series of the Spanish Day-Ahead market.
4.3.1 Data
Data for the Day-Ahead market price in Spain and 3 different one day-ahead point forecasters ($M=3$, Figure 8(a)) is available from 1st of January 2023 to 28th of December 2024. The period from 1st of January 2024 to 28th of December 2024 is considered as test data. This is the window where conformal methods are applied. In order for these methods to be applied, quantile regression forecasts must be available. These will be obtained from 5th of July 2023 (Figure 8(b)). This period is ideal for testing this type of models because moments of great uncertainty can be observed at the same time as very steady phases. The point forecasting models are related to the ones presented in Lago et al. (2021); Sebastián et al. (2023) [^3] They can’t be completely described for commercial reasons..
4.3.2 Methodology
When forecasting the price one day in advance, it must be predicted for the 24 hours of the next day. Thus, there are two ways of proceeding with the quantile regression: consider 24 daily series (one per hour) and a quantile regression model for each one of these series or a single hourly frequency series and therefore a single quantile regression model for all hours. As not many data is available, only the second approximation is considered. In any case, both approximations are valid and it has been shown that, for point forecasting scenarios, results are quite similar (Ziel and Weron, 2018). A training rolling window of 180 days is considered. That is, at the time of predicting the day $D$, data from days $D-180$ to $D-1$ are considered to train the quantile regression models. For the day $D+1$, the window from $D-179$ to $D$ is considered for training and so on. For those days for which the training window is shorter than 180 days due to data unavailability, all available data will be used. Conformalization is always carried out individually for each hour. For both ACI and WACI $\gamma=0.02$ is taken, which seems reasonable in view of previous studies (Zaffran et al., 2022) and for the WACI approach $\sigma=3$, which also seems appropriate given the price scale. For WACI, a grid $\bm{L}$ based on a 0.1€/MWh step is used. With all these distinctions, 9 possible methodologies will be compared over a test period of almost one year. It is crucial to include at least one year of data to ensure the generalizability of the results (Lago et al., 2021). A full year encompasses special situations such as holidays, demand seasonality, and varying meteorological conditions. This approach prevents the results from being biased or influenced by the exclusion of any particular situation, providing a comprehensive evaluation of the methods. Two possible values of $\alpha$ will be distinguished: $0.2,0.1$, which correspond to coverage values of 80, 90%, respectively. The calibration window size, as well as the $\gamma$ and $\sigma$ values, are hyperparameters that could be optimized to improve performance. However, in this study, such optimization has not been conducted, indicating that better results may be achievable with further tuning. To better understand the impact of $\sigma$ on the WACI methodology and the differences with ACI, an analysis can be found in the B. In this example, estimations of the standard deviation for each metric have been obtained through the stationary bootstrap procedure from Politis and Romano (1994), considering 1000 bootstrapped samples of size 1000.
4.3.3 Results and discussion
Results for every evaluation metric are shown in Tables 4 and 5, for $\alpha=0.20$ and $\alpha=0.10$, respectively. Metric names are displayed with their abbreviations, for better readability of the tables. On the left side of the tables, the standard metrics of mean empirical coverage, mean interval length and Winkler score are displayed. On the right side, conditional coverage metrics as well as those related to desired properties are shown.
Methodology M.E.C. M.I.L. W.S. P.C. ILS 0.10 S.C. I.L. Std MCD 5
QRA
79.68
32.07
49.85
0.16
0.10
9.43
7.43
HQR
80.24
31.15
47.48
0.04
0.32
12.54
4.83
HQR-W
80.78
30.95
47.11
0.05
0.30
12.04
5.55
QRA (ACI)
79.20
32.20
49.13
0.20
0.60
0.20
12.27
6.91
HQR (ACI)
79.62
31.77
47.13
0.12
0.40
0.35
14.86
4.86
HQR-W (ACI)
79.50
31.21
46.81
0.14
0.21
0.33
14.17
5.69
QRA (WACI)
80.67
32.33
48.99
0.14
1.15
0.16
10.03
5.38
HQR (WACI)
79.90
31.16
47.09
0.08
0.08
0.33
13.20
3.84
HQR-W (WACI)
80.25
31.17
46.80
0.08
0.09
0.31
12.52
4.42
Table 4: Evaluation metrics for the EPF example for $\alpha=0.20$. Standard deviations of each metric are shown in brackets.
Methodology M.E.C. M.I.L. W.S. P.C. ILS 0.10 S.C. I.L. Std MCD 5
QRA
89.10
44.86
63.62
0.14
0.08
11.45
4.65
HQR
89.58
43.11
59.67
0.06
0.31
15.79
3.64
HQR-W
89.61
42.45
59.75
0.06
0.29
14.36
3.25
QRA (ACI)
89.50
46.36
63.04
0.16
0.41
0.18
16.50
4.54
HQR (ACI)
89.84
43.81
59.38
0.11
0.08
0.34
19.53
3.32
HQR-W (ACI)
89.61
44.04
59.59
0.11
0.24
0.31
18.78
3.80
QRA (WACI)
90.31
46.65
62.98
0.13
0.38
0.13
12.97
3.72
HQR (WACI)
90.14
43.52
59.35
0.07
0.06
0.32
17.60
2.57
HQR-W (WACI)
90.03
43.90
59.65
0.07
0.13
0.30
16.90
2.66
Table 5: Evaluation metrics for the EPF example for $\alpha=0.10$. Standard deviations of each metric are shown in brackets.
In order to have a better understanding of the base behaviour of the models, a plot has been made of the mean empirical coverage against the mean interval length (Figure 9).
Figure 9: Mean empirical coverage vs. mean interval length for both levels of $\alpha$, 0.10 and 0.20. The colour of each point is determined by the base quantile regression model used. The shape of each point is determined by the applied conformalization. The WACI-HQR combined methodology is highlighted with a red circle.
Comparison between quantile regression models
At first glance, it is evident that models based on QRA perform significantly worse than those based on HQR(-W), regardless of the chosen $\alpha$ level. This is reflected in the Winkler score, where the difference is particularly pronounced, indicating poorer overall performance for QRA. Delving deeper into the metrics, several shortcomings of QRA become apparent. First, the Pearson correlation for QRA is notably higher, signalling a weaker level of conditional coverage compared to HQR(-W). Additionally, the relationship between interval length and prediction difficulty, as measured by the Spearman correlation, is considerably weaker for QRA. This suggests that QRA struggles to adjust interval lengths appropriately to reflect varying levels of uncertainty. Further evidence of this shortcoming is found in the lower standard deviation of interval lengths for QRA. This indicates an inability to differentiate between contexts of varying prediction difficulty, which complicates decision-making processes. Finally, while none of the methods are explicitly designed for this purpose, the superior approximation of quantiles achieved by HQR(-W) also benefits the MCD metric. In conclusion, QRA fails to match the quality achieved by HQR(-W) across all evaluated metrics.
Inclusion (or not) of varying weights in the HQR model
The differences between HQR and HQR-W are minimal and not statistically significant, suggesting that both models exhibit nearly identical behaviour. This observation holds true even when analysing different conformalization methods applied to the models. Based on the principle of parsimony, this suggests that HQR should be preferred. By selecting the simpler model, one can avoid potential overfitting issues that might arise from the additional complexity introduced by HQR-W.
Differences between ACI and WACI
When examining traditional metrics, ACI and WACI produce very similar results and are nearly indistinguishable, though two key observations stand out. First, WACI consistently achieves better Winkler score values when comparing the same base models and, second, produces valid prediction intervals, except for the HQR (WACI) case with $\alpha=0.20$ by a very small and reasonable margin. Pearson’s correlation reveals a clear difference: when applied to the QRA model, WACI improves conditional coverage, whereas ACI worsens it. For the HQR(-W) base model, ACI considerably degrades this property (with the coefficient being over 0.1), while WACI maintains it at levels comparable to the original model. Further insights emerge from the ILS 0.10 metric, where WACI demonstrates clear superiority, achieving minimal values and bringing interval coverage closer to the target. This robustness gives users confidence, as the methodology performs well even with significant ex-post corrections. Regarding the Spearman correlation and the standard deviation of interval lengths, WACI and ACI preserve these characteristics from the base models, a positive outcome that aligns with the intent behind using HQR. Finally, on the MCD metric, WACI once again outperforms ACI with a notable difference, demonstrating greater independence between the coverage indicator and interval length, which is the idea behind WACI’s design. In conclusion, while ACI and WACI often appear similar in their outcomes, WACI consistently demonstrates superior performance across desirable properties.